# MRGen: Diffusion-based Controllable Data Engine for MRI Segmentation towards Unannotated Modalities

### v1

@@ Appendix D More Experiments
In this section, we present additional experimental results to demonstrate the superiority of our proposed data engine. In Section D.1, we showcase both quantitative and qualitative results of in-domain generation. In Section D.2, we conduct extensive ablation studies to illustrate the improvements introduced by our design in radiology image generation. Finally, in Section D.3, we provide additional qualitative results to validate the accuracy and flexibility of the generated outputs.
@@ D.1 In-domain Generation
Our proposed data engine not only synthesizes images for target modalities lacking mask annotations but also maintains controllable generation capabilities within the source domain. Moreover, as presented in Table 6, downstream segmentation models trained exclusively on synthetic source-domain data can achieve performance comparable to those trained on real manually-annotated source-domain data. This offers a feasible solution to address concerns about medical data privacy.
@@
Table 6: More Quantitative Results (DSC score) on Segmentation. We compare the performance of nnUNet [21] trained on real data versus synthetic data generated by our MRGen in both the source domain ($\mathcal{D}_{s}$) and target domain ($\mathcal{D}_{t}$).
@@
Figure 7: Qualitative Results of In-domain Generation.
D.2 Ablation Studies
In this section, we provide more ablation study results on image generation, to demonstrate the necessity of model and data design within our proposed data engine, focusing on autoencoder reconstruction and text-guided generation.
Autoencoder reconstruction. We evaluate the reconstruction performance of various VAEs on the test set derived from our curated MedGen-1M dataset, comprising Radiopaedia-CT and Radiopaedia-MRI. The evaluated models include: (i) the VAE of pretrained Stable Diffusion (VAE-SDM), (ii) the VAE of Stable Diffusion finetuned on our dataset (Finetuned-VAE-SDM), (iii) our model trained solely on Radiopaedia-MRI data (VAE-MRGen (MRI-only)), and (iv) our VAE-MRGen trained on the complete dataset (VAE-MRGen (Ours)).
The reconstruction quality is assessed using standard metrics, including Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), L1 Error, and Mean Squared Error (MSE) between the reconstructed and original images.
As illustrated in Table 8, we draw the following three key observations: (i) Finetuning on radiology data: VAEs finetuned on radiology data significantly improve the reconstruction quality compared to pretrained Stable Diffusion models, revealing the limited generalizability of natural image VAEs to radiology images. (ii) Latent code dimension: our autoencoders with a larger latent code feature dimension achieve further improvements in reconstruction quality, suggesting that increasing the increased capacity of latent code enables better representations. (iii) Extra data: incorporating CT data into training further improves reconstruction quality, approaching the performance of lossless compression. To this end, unlike RoentGen [3] that relies on pretrained VAEs, we employ high-capacity autoencoders trained specifically on radiology data, to ensure that the generative models built upon this foundation can produce better image details.
Text-guided generation. We evaluate the text-guided image generation performance of various models on the test set derived from our MedGen-1M dataset, encompassing Radiopaedia-CT and Radiopaedia-MRI. The evaluated models include: (i) the pretrained Stable Diffusion (SDM), (ii) the Stable Diffusion finetuned on our MedGen-1M (Finetuned-SDM), (iii) our model using only free-text modality labels as text conditions (MRGen (Modality-only)), and (iv) our MRGen, which utilizes templated text prompts incorporating modality labels, attributes, regions, and organs information (MRGen (Ours)).
We employ commonly used evaluation metrics for image generation, including the Fréchet Inception Distance (FID [16]) to assess the quality and diversity of generated images. Additionally, we utilize BiomedCLIP [52] to calculate two similarity metrics: image-to-image similarity (CLIP-I) between the generated images and ground truth, and image-to-text similarity (CLIP-T) between the generated images and the free-text modality descriptions.
As presented in Table 8, we draw the following three observations: (i) Finetuning on radiology data: pretrained Stable Diffusion tends to generate natural images rather than ideal radiology images, resulting in suboptimal performance across all evaluation metrics. However, finetuning on our dataset significantly enhances its generative performance. (ii) Text encoder: using BiomedCLIP [52] as the text encoder, trained on biomedical data, instead of CLIP [40] in Stable Diffusion, improves MRGen’s ability to capture similarities and differences among radiology modalities. Additionally, our high-capacity VAE, provides further potential improvements to generation quality. (iii) Templated text prompt: incorporating templated text prompts with detailed information (including modality, attributes, regions, and organs) outperforms using modality labels alone. This enables MRGen to better grasp the relationships among distinct modalities, thereby further improving generation quality, particularly in terms of image-to-image similarity with the ground truth.
Model Dim PSNR $\uparrow$ SSIM $\uparrow$ L1 Error $\downarrow$ MSE $\downarrow$
VAE-SDM [42] 4 31.32 0.989 0.0294 0.0037
Finetuned-VAE-SDM 4 35.65 0.996 0.0190 0.0014
VAE-MRGen (MRI-only) 16 40.56 0.998 0.0115 0.0005
VAE-MRGen (Ours) 16 42.62 0.999 0.0091 0.0003
Table 7: Ablation Studies on Autoencoder Reconstruction.
Model FID $\downarrow$ CLIP-I $\uparrow$ CLIP-T $\uparrow$
SDM [42] 249.24 0.3151 0.1748
Finetuned-SDM 91.48 0.6698 0.3199
MRGen (Modality-only) 41.82 0.7512 0.3765
MRGen (Ours) 39.63 0.8457 0.3777
Table 8: Ablation Studies on Text-guided Generation.
@@ D.3 More Qualitative Results
Image generation. As depicted in Figure 9 and 10, we present additional visualizations of controllable generation on target modalities lacking mask annotations. These results demonstrate that the proposed data engine can effectively generate high-quality samples based on masks across various datasets and modalities, facilitating the training of downstream segmentation models towards these challenging scenarios.
Image segmentation. As presented in Figure 11 and 12, we provide more visualizations of segmentation models trained using synthetic data on modalities that originally lack mask annotations. This validates that the samples generated by MRGen can effectively assist in training segmentation models, achieving impressive performance in previously unannotated scenarios.
@@ E.2 Future Works
To address the aforementioned limitations of MRGen, we propose several directions for future improvement: (i) Constructing more comprehensive and richly annotated datasets, such as incorporating more annotated CT data or collecting more MRI data, to enhance the model’s ability to effectively utilize mask conditions; (ii) Designing finer-grained, accurate, and efficient generative model architectures to improve generation efficiency and accuracy, particularly for small-volume organs; and (iii) Developing a more robust and comprehensive data filtering pipeline to reliably select high-quality samples that meet the requirements of downstream tasks.
Figure 9: More Qualitative Results of Controllable Generation. We present images from the source domain $\mathcal{D}_{s}$ and the target domain $\mathcal{D}_{t}$ for reference. Here, specific organs are contoured with distinct colors: prostate in MSD-Prostate and PROMISE12 datasets, and pancreas in PanSeg dataset, and liver, right kidney, left kidney and spleen in other datasets.
Figure 10: More Qualitative Results of Controllable Generation. We present images from the source domain $\mathcal{D}_{s}$ and the target domain $\mathcal{D}_{t}$ for reference. Here, specific organs are contoured with distinct colors: prostate in MSD-Prostate and PROMISE12 datasets, and pancreas in PanSeg dataset, and liver, right kidney, left kidney and spleen in other datasets.
Figure 11: More Qualitative Results on Segmentation towards Unannotated Modalities. We present reference images from the source-domain training set $\mathcal{D}_{s}$, images from the target-domain test set $\mathcal{D}_{t}$, and corresponding predictions and ground truth. Here, specific organs are highlighted with different colors: prostate in MSD-Prostate and PROMISE12 datasets, and pancreas in PanSeg dataset, and liver, right kidney, left kidney and spleen in other datasets.
Figure 12: More Qualitative Results on Segmentation towards Unannotated Modalities. We present reference images from the source-domain training set $\mathcal{D}_{s}$, images from the target-domain test set $\mathcal{D}_{t}$, and corresponding predictions and ground truth. Here, specific organs are highlighted with different colors: prostate in MSD-Prostate and PROMISE12 datasets, and pancreas in PanSeg dataset, and liver, right kidney, left kidney and spleen in other datasets.

### v2

@@ Appendix D More Experiments
In this section, we present additional experimental results to demonstrate the superiority of our proposed data engine. First, in Sec. D.1, we showcase quantitative and qualitative results of in-domain generation. Next, in Sec. D.2, we present quantitative comparisons with more baselines, further confirming the effectiveness and necessity of our proposed data engine. Finally, in Sec. D.3, we provide extra qualitative results to validate the accuracy and flexibility of the generated outputs.
@@ D.1 In-domain Generation
Our proposed data engine not only synthesizes images for target modalities lacking mask annotations but also maintains controllable generation capabilities within the source domains. Moreover, as presented in Table 8, downstream segmentation models trained exclusively on synthetic source-domain data can achieve performance comparable to those trained on real manually-annotated data. This offers a feasible solution to address concerns about medical data privacy.
@@
Table 8: In-domain & Cross-domain Augmentation Results (DSC score) on Segmentation. We compare the performance of nnUNet [23] trained on real data versus synthetic data generated by MRGen in both the source domain ($\mathcal{D}_{s}$) and target domain ($\mathcal{D}_{t}$).
Figure 7: Qualitative Results of In-domain Generation.
@@
Dataset Source Modality Target Modality DualNorm nnUNet
$\mathcal{D}_{s}$ MRGen CycleGAN UNSB MaskGAN MAISI Few-shot UniVerSeg TS-MRI Oracle Box SAM2 nnUNet $\mathcal{D}_{t}$
CHAOS-MRI T1 T2-SPIR 14.00 6.90 66.18 7.58 14.03 32.73 3.34 52.00 48.91 80.64 45.45 53.12 83.90
T2-SPIR T1 12.50 0.80 58.10 1.38 6.44 1.89 3.11 53.82 52.79 77.09 43.48 51.94 90.60
MSD-Prostate T2 ADC 1.43 5.52 57.83 40.92 52.99 29.14 9.15 20.28 0.0 0.0 61.50 65.39 82.35
ADC T2 12.94 22.20 61.95 57.06 38.39 5.98 6.94 29.38 53.90 0.0 61.07 66.40 89.80
Average DSC score 10.22 8.86 61.02 26.74 27.96 17.44 5.64 38.87 38.90 39.43 52.88 59.21 86.66
Table 9: More Quantitative Results (DSC score) on Segmentation. The best and second-best performances are bolded and underlined, respectively. Notably, the results marked with a gray background indicate that the corresponding methods may have accessed target-modality annotated data during extensive training (e.g., UniVerSeg, TotalSegmentor-MRI (TS-MRI)), utilized oracle inputs as prompts (e.g., Oracle Box, SAM2), or even been directly trained on target-modality annotated data (e.g., nnUNet (Few-shot), nnUNet ($\mathcal{D}_{t}$)). Consequently, these approaches do not represent a fully fair comparison with others, and are primarily included as performance references.
D.2 More Quantitative Results
In this section, we compare MRGen with additional baseline methods on two typical cross-modal dataset pairs from MRGen-DB by evaluating the performance of downstream segmentation models, as detailed in the main text. Concretely, for both translation-based and generation-based methods, we assess the performance of nnUNet [23] trained on data generated by these methods. As depicted in Table 9, we further analyze the relevant baselines by category, as follows.
Augmentation-based methods. Limited to relying on carefully crafted augmentation strategies, DualNorm [68] fails to model nonlinear visual discrepancies among distinct modalities, leading to poor cross-modality segmentation performance.
Translation-based methods. While CycleGAN [69], UNSB [27], and MaskGAN [46] excel at contour preservation, they often suffer from model collapse when learning complex modality transformations, resulting in suboptimal performance.
Generation-based models. Despite finetuned on our dataset, the performance of MAISI [14] is still poor, which we attribute to its lack of modality-conditioning capability. This limitation hinders its ability to support cross-modality generation, and consequently, makes it struggle to synthesize target-domain samples for training segmentation models.
Few-shot methods. While few-shot nnUNet [23] and UniverSeg [5] benefit from partial target-domain annotations, MRGen-boosted models outperform without requiring any such annotations, showcasing practical feasibility in clinical scenarios.
General segmentation models. TotalSegmentor-MRI [8] works well on certain datasets/modalities (likely already included during training), but it still performs poorly or even fails on others. This significantly limits its practicality in complex clinical scenarios, especially when dealing with underrepresented modalities with diverse imaging characteristics.
Models with oracle inputs. Although SAM2 [48] with perturbed oracle boxes as prompts exhibits impressive zero-shot segmentation capabilities, our MRGen-boosted models still outperform it, trailing only the oracle nnUNet trained directly on target-domain annotated data. Moreover, as a semi-automatic method, SAM2’s reliance on high-quality spatial prompts and manual intervention limits its scalability and applicability, while MRGen offers a fully automated, end-to-end solution.
Overall, MRGen provides a robust, fully automated approach for challenging cross-modality segmentation by producing high-quality synthetic data, with no need for any target-domain mask annotations and proving highly suitable for clinical applications. For computational efficiency, we primarily focus on comparing MRGen with some representative baselines, DualNorm [68], CycleGAN [69] and UNSB [27], across more dataset pairs in the main text for a comprehensive evaluation.
@@ D.3 More Qualitative Results
Image generation. As depicted in Figure 9, we present additional visualizations of controllable generation on target modalities lacking mask annotations. These results demonstrate that the proposed data engine can effectively generate high-quality samples based on masks across various datasets and modalities, facilitating the training of downstream segmentation models towards these challenging scenarios.
Image segmentation. As presented in Figure 10, we provide more visualizations of segmentation models trained using synthetic data on modalities that originally lack mask annotations. This validates that the samples generated by MRGen can effectively assist in training segmentation models, achieving impressive performance in previously unannotated scenarios.
@@ E.2 Future Works
Due to limited computational resources, we validate our proposed data engine on 2D slices in this paper, with trained segmentation models able to process 3D volumes slice-by-slice. However, we believe our idea can be seamlessly extended to 3D volume generation. With more computing in the future, we aim to develop a 3D version model for the community, further advancing cross-modality segmentation performance. Moreover, to address the aforementioned limitations of MRGen, we propose several directions for future improvement: (i) Constructing more comprehensive and richly annotated datasets, such as incorporating more annotated MRI data, to enhance the model’s ability to effectively utilize mask conditions; (ii) Designing finer-grained, accurate, and efficient generative model architectures to improve generation efficiency and accuracy, particularly for small-volume organs; and (iii) Developing a more robust and comprehensive data filtering pipeline to reliably select high-quality samples that meet the requirements of downstream tasks.
Figure 9: More Qualitative Results of Controllable Generation. We present images from source domains $\mathcal{D}_{s}$ and target domains $\mathcal{D}_{t}$ for reference. Here, specific organs are contoured with colors: prostate in MSD-Prostate and PROMISE12 datasets, and pancreas in PanSeg dataset, and liver in LiQA and CHAOS-MRI datasets.
Figure 10: More Qualitative Results on Segmentation towards Unannotated Modalities. Significant imaging differences between source-domain ($\mathcal{D}_{s}$) and target-domain ($\mathcal{D}_{t}$) make segmentation on target domains ($\mathcal{D}_{t}$) extremely challenging. Here, specific organs are highlighted with colors: prostate in MSD-Prostate and PROMISE12, pancreas in PanSeg, and liver in LiQA and CHAOS-MRI datasets.